# YOLO11 to Its Genesis: A Decadal and Comprehensive Review of The You Only Look Once (YOLO) Series

### v5

@@ IV-A Challenges in Statistical Metrics for Evaluation
Threat: Relying on a single statistical summary metric to measure YOLO detection capability may not fully reflect the performance of systems across various YOLO applications, necessitating the use of several metrics.
@@
V Emerging Trends and Future Directions in Object Detection with YOLO
In considering future challenges, it is envisioned that YOLO variants will continue to improve performance on small object targets, especially as they penetrate into more specialized areas such as precision manufacturing and robotic fruit harvesting. Attention mechanisms could be integrated to enhance small object detection, while the use of vision transformers could further improve YOLO’s performance on capturing global contextual dependencies.
As YOLO progresses, meeting the demands of niche applications will likely drive further innovation in architectural design and optimization, ensuring its continued relevance in domains with stringent requirements for precision and efficiency. For instance, the integration of YOLO and voice commands into surveillance systems, facilitated by models like generative pretrained transformer, could transform security mechanisms, making them more interactive and responsive. In the healthcare sector, incorporating medical imaging with historical patient data and live symptom descriptions could significantly improve the personalization and accuracy of medical responses. In autonomous vehicles, YOLO could be utilized to detect objects like road signs and pedestrians, enabling the planning of safe and efficient vehicle navigation. In agriculture, the integration of YOLO and language models would enhance the interactive diagnosis of crop stresses and provide recommendations for specific stress management.
Looking further, YOLO’s potential to adapt to such multimodal advancements will be instrumental in pioneering the next wave of intelligent applications. This transformative phase promises to significantly improve the interactivity and cognitive capabilities of Machine Vision systems, marking a pivotal shift in visual process automation.
V-A YOLO and Artificial General Intelligence
Artificial General Intelligence (AGI) refers to an intelligent agent with human-level or higher intelligence, capable of solving a variety of complex problems in diverse domains [189, 190]. In this context, an AGI system would need to integrate object detection capabilities, akin to those provided by YOLO, with other essential cognitive functions, such as advanced natural language understanding, reasoning, and decision-making. This fusion would enable the system to handle a broad spectrum of tasks in real time, adapting to dynamic environments and complex scenarios effectively.
V-B YOLO Integration with Large Language Models
Integrating YOLO with Large Language Models (LLMs) could substantially elevate an AGI’s capabilities by seamlessly merging advanced visual data interpretation with sophisticated natural language understanding, reasoning, and contextual awareness. This fusion would allow the AGI to not only recognize and analyze objects in real-time but also engage in meaningful interactions, making more informed decisions and adapting to complex tasks with greater autonomy and precision. This synergy would enable AGI systems to operate across multiple fronts simultaneously, such as navigating through an environment using visual cues identified by YOLO while interpreting and acting on spoken commands through capabilities provided by LLMs [191, 192, 128]. Such integration seeks to develop a highly versatile and intelligent system, capable of performing real-time, multi-faceted operations across diverse domains. By combining advanced visual recognition with robust language processing, it brings us a step closer to realizing true AGI, where the system can autonomously adapt, learn, and perform complex tasks with human-like flexibility and reasoning.
V-C YOLO deployment on Edge Devices
@@
The deployment of YOLO on edge devices unlocks several promising avenues for future research and development. One potential direction involves enhancing the algorithm’s efficiency and accuracy for even more constrained environments, such as ultra-low-power Microcontrollers and embedded systems. This can be achieved through further optimization techniques, including model pruning, quantization, and the development of specialized hardware accelerators. Additionally, integrating YOLO with advanced communication protocols and edge computing frameworks could facilitate more seamless collaboration between edge devices and centralized cloud services, enhancing the overall system performance and scalability. Exploring the integration of YOLO with other AI-driven functionalities, such as anomaly detection and predictive analytics, may unlock new applications in areas like healthcare, smart cities, and industrial automation. As edge computing continues to evolve, the adaptation of YOLO to support federated learning paradigms could ensure the data privacy while enabling continuous learning and improvement of object detection models. These future directions will not only expand the capabilities of YOLO but also contribute significantly to the advancement of intelligent edge computing systems [193, 194, 195].
V-D YOLO and Embodied Artificial Intelligence
@@
V-E YOLO and Environmental Impact
@@
VI Conclusion
In this comprehensive review, we explored the evolution of the YOLO models from the most recent YOLOv11 to the inaugural YOLOv1. Our review outlined the significant enhancements in detection speed, accuracy, and computational efficiency that each iteration brought, while also addressing the specific challenges and limitations faced by earlier versions. Furthermore, we identified gaps in the current capabilities of YOLO models and proposed potential directions for future research, such as trade-off between detection speed versus accuracy, handling small and overlapping Objects, and generalization across diverse datasets and domains. Predicting the trajectory of YOLO’s development, we anticipate a shift towards multimodal data processing, leveraging advancements in large language models and natural language processing to enhance object detection systems. This fusion is expected to broaden the utility of YOLO models, enabling more sophisticated, context-aware applications that could revolutionize interaction between AI systems and their environments using Generative AI and multi-modal LLMs. Thus, this review not only serves as a detailed chronicle of YOLO’s evolution but also sets a prospective blueprint for its integration into the next generation of technological innovations.
VII Authors Contribution
Ranjan Sapkota: principal conceptualizer, research design, formal analysis, original draft preparation, manuscript writing, and editing. Rizwan Qureshi, Marco Flores-Calero, Chetan Badgujar, Upesh Nepal, Alwin Poulose, Peter Zeno, Uday Bhanu Prakash Vaddevolu, Sheheryar Khan, Maged Shoman, Hong Yan: methodology refinement, critical revisions, manuscript review, and editing. Manoj Karkee: Funding and Supervision, methodology refinement, critical revisions, manuscript review, and editing. Ranjan Sapkota and Manoj Karkee: Corresponding authors.
VIII Acknowledgement
This work was supported by the National Science Foundation and United States Department of Agriculture, National Institute of Food and Agriculture through the “Artificial Intelligence (AI) Institute for Agriculture” Program under Award AWD003473. Additionally, this work is supported in parts by Hong Kong Innovation and Technology Commission (InnoHK Project CIMDA).
Authors’ Contribution: Ranjan Sapkota led research design, analysis, and manuscript drafting. Rizwan Qureshi, Syed Zohaib Hassan, John Shutske, Maged Shoman, Muhammad Sajjad, Fayaz Ali Dharejo, Achyut Paudel, Jiajia Li, Zhichao Meng, Ferhat Sadak, and Muhammad Usman Hadi refined methods and reviewed critically. Manoj Karkee facilitated funding and manuscript drafting/edit. Supported by NSF and USDA under the ’AI Institute for Agriculture’ program, Award AWD003473.

### v6

@@ IV-A Challenges in Statistical Metrics for Evaluation
Threat: Evaluating YOLO detection systems requires a unique approach, as each version, from the original YOLO to the latest YOLOv12, targets different aspects of detection capability, such as speed, accuracy, or computational efficiency. For a comprehensive evaluation, it is essential to employ a diverse array of metrics, including precision, recall, GFLOPs, and model size. This approach allows for a more complete comparison and understanding of each model’s strengths and weaknesses in various real-world applications. This multi-metric evaluation is crucial to assessing the practical utility and technological advancement of the YOLO series. Future YOLO versions are expected to introduce novel evaluation metrics that capture emerging capabilities in edge computing, multi-modal fusion, and adaptive architecture optimization, necessitating an even more sophisticated evaluation framework to accurately assess their performance across diverse deployment scenarios.
@@
V Future Directions in Object Detection with YOLO
V-A YOLO deployment on Edge and IoT Devices
@@
The deployment of YOLO on edge devices unlocks several promising avenues for future research and development. One potential direction involves enhancing the algorithm’s efficiency and accuracy for even more constrained environments, such as ultra-low-power Microcontrollers and embedded systems. This can be achieved through further optimization techniques, including model pruning, quantization, and the development of specialized hardware accelerators. Additionally, integrating YOLO with advanced communication protocols, edge computing frameworks and IoT devices could facilitate more seamless collaboration between edge devices and centralized cloud services, enhancing the overall system performance and scalability. Exploring the integration of YOLO with other AI-driven functionalities, such as anomaly detection and predictive analytics, may unlock new applications in areas like healthcare, smart cities, and industrial automation. As edge computing continues to evolve, the adaptation of YOLO to support federated learning paradigms could ensure the data privacy while enabling continuous learning and improvement of object detection models. These future directions will not only expand the capabilities of YOLO but also contribute significantly to the advancement of intelligent edge computing systems [174, 175, 176, 177].
V-B YOLO and Embodied Artificial Intelligence
@@
VI Expanding YOLO Object Detection into broader AI Domains
VI-A YOLO and Artificial General Intelligence
Artificial General Intelligence (AGI) refers to an intelligent agent with human-level or higher intelligence, capable of solving a variety of complex problems in diverse domains [183, 184]. In this context, an AGI system would need to integrate object detection capabilities, similar to those provided by YOLO, with other essential cognitive functions, such as advanced natural language understanding, reasoning, and decision-making. This fusion will enable the system to handle a broad spectrum of tasks in real-time, adapting to dynamic environments and complex scenarios effectively, thus advancing the current AI systems towards achieving a true AGI.
VI-B YOLO Integration with Large Language Models
One effective way to advance AGI capabilities is to integrate YOLO with Large Language Models (LLMs) by seamlessly merging advanced visual data interpretation with sophisticated natural language understanding, reasoning, and contextual awareness. This fusion would allow the AGI to not only recognize and analyze objects in real-time but also engage in meaningful interactions with stakeholders (or end users), making more informed decisions and adapting to complex tasks with greater autonomy and precision. This synergy would enable AGI systems to operate in complex environments handling multi-modal inputs simultaneously, such as navigating through an environment using visual cues identified by YOLO while interpreting and acting on spoken commands through capabilities provided by LLMs [185, 186, 112]. Such integration is expected to lead to a highly versatile and intelligent system, capable of performing real-time, multi-faceted operations across diverse application domains. By combining advanced visual recognition with robust language processing, it brings us a step closer to realizing true AGI, where the system can autonomously adapt, learn, and perform complex tasks with human-like flexibility and reasoning.
VII YOLO and Environmental Impact
@@
VIII Conclusion
In this comprehensive review, we explored the evolution of the YOLO models from the most recent YOLOv12 to the inaugural YOLOv1, including alternative versions of YOLO as YOLO-NAS, YOLO-X, YOLO-R, DAMO-YOLO, and Gold-YOLO. This retrospective analysis covered a decade of advancements, highlighting theapplied use of each version and their respective impacts across five critical application areas: autonomous vehicles and traffic safety, healthcare and medical imaging, security and surveillance, manufacturing, and agriculture. Our review outlined the significant enhancements in detection speed, accuracy, and computational efficiency that each iteration brought, while also addressing the specific challenges and limitations faced by earlier versions. Furthermore, we identified gaps in the current capabilities of YOLO models and proposed potential directions for future research, such as trade-off between detection speed versus accuracy, handling small and overlapping Objects, and generalization across diverse datasets and domains. Predicting the trajectory of YOLO’s development, we anticipate a shift towards multimodal data processing, leveraging advancements in large language models and natural language processing to enhance object detection systems. This fusion is expected to broaden the utility of YOLO models, enabling more sophisticated, context-aware applications that could revolutionize the interaction between AI systems and their environments using Generative AI and multi-modal LLMs. Thus, this review not only serves as a detailed chronicle of YOLO’s evolution but also sets a prospective blueprint for its integration into the next generation of technological innovations.
IX Authors Contribution
Ranjan Sapkota: principal conceptualizer, research design, formal analysis, original draft preparation, manuscript writing, and editing. Marco Flores-Calero, Rizwan Qureshi, Chetan Badgujar, Upesh Nepal, Alwin Poulose, Peter Zeno, Uday Bhanu Prakash Vaddevolu, Sheheryar Khan, Maged Shoman, Hong Yan: methodology refinement, critical revisions, manuscript review, and editing. Manoj Karkee: Funding and supervision, methodology refinement, critical revisions, manuscript review, and editing. Ranjan Sapkota and Manoj Karkee: Corresponding authors.
X Acknowledgement
This work was supported by the National Science Foundation and United States Department of Agriculture, National Institute of Food and Agriculture through the “Artificial Intelligence (AI) Institute for Agriculture” Program under Award AWD003473. Additionally, this work is partially supported by the Hong Kong Innovation and Technology Commission (InnoHK Project CIMDA) and by EcuTS2025 from UFA-ESPE.
Our more research on CV and YOLO [79], [189], [112], [190], [191], [192], [193], [194], [195], [196]